Ab Initio Vibrational Splittings for Hydrogen Atom Transfer in HO_{2}^{−}/O_{2}H^{−} Isomerization

\underline{Wai-To Chan} and Ian P. Hamilton, Department of Chemistry, Wilfrid Laurier University, Waterloo, Ontario..

A potential energy surface covering the interconversion between the two symmetry-equivalent geometries of the hydroperoxyl anion, HO_{2}^{−} via a symmetric T-shaped saddle point was computed at the QCISD(T)/6-311++G(2df,pd) level of theory. An analytical molecular potential function was fit to the ab initio data. The classical barrier to the interconversion was estimated to be 6057 cm^{−1} Using the parameterized function the vibrational states were calculated by the discrete variable representation-distributed Gaussian basis method. Computations predict the onset of tunnelling from about 4 quanta in the bending mode-A splitting of 6 cm^{−1} is observed between the odd and even 040 state. Presence of at least one quanta in the O–O stretching mode enhances splitting significantly.